# Efficient Fairness-Performance Pareto Front Computation

### v1

@@ 4 The Invertibility Theorem
In this section we define the notion of invertibility for representations, and show that optimal fair representations are invertible.
To this end, let us first introduce an additional notation, which is more convenient when describing representations specifically on finite sets. Let $S_{0},S_{1}$ be finite disjoint sets, where $S_{a}$ represents the values of $(X,A)$ when $A=a$, for $a\in\left\{0,1\right\}$. Denote $S=S_{0}\cup S_{1}$. We are assuming that there is a probability distribution $\zeta$ on $S$, and $A$ is the random variable $A=\mathbbm{1}_{\left\{s\in S_{1}\right\}}$. $X$ is defined as taking the values $s\in S$, with $\mathbb{P}\left(X=s\right)=\zeta(s)$. Further, the variable $Y$ is defined to take values in a finite set $\mathcal{Y}$, and for every $s\in S$, its conditional distribution is given by $\rho_{s}\in\Delta_{\mathcal{Y}}$. That is, $\mathbb{P}\left(Y=y|X=s,A=a\right)=\rho_{s}(y)=\rho_{s,a}(y)$.  [^2] Note that there is a slight redundancy in the notation $\mathbb{P}\left(Y=y|X=s,A=a\right)$ here, since $a$ is determined by $s$. However, to retain compatibility with the standard notaion, we specify them both. This is similar to the continuous situation, in which although $A$ is technically part of the features, $X$ and $A$ are specified separately. This completes the description of the data model.
We now describe the representation. The representation will take values in a finite set $\mathcal{Z}$. For every $s\in S$ and $z\in\mathcal{Z}$, let $T_{a}(z,s)=\mathbb{P}\left(Z=z|X=s,A=a\right)$ be the conditional probability of representing $s$ as $z$. Then $T_{a}$’s fully define the distribution of the representation $Z$. We shall refer to the representation as $T$ or $Z$ in interchangeably. Finally, for $a\in\left\{0,1\right\}$ denote
Similarly to the previous section, our goal is to find representations $T$ that minimise the cost (5) under the constraint (3), for some fixed $\gamma>0$.
@@
A representation $T$ is invertible if for every $z\in\mathcal{Z}$ and every $a\in\left\{0,1\right\}$, there is at most one $s\in S_{a}$ such that $T_{a}(z,s)>0$. In words, a representation is invertible, if any given $z$ can be produced by at most two original features $s$, and at most one in each of $S_{0}$ and $S_{1}$.
@@
Let $T$ be any representation of $(S,A,\zeta,Y)$ on a set $\mathcal{Z}$. Then there exists an invertible representation $T^{\prime}$ of $(S,A,\zeta,Y)$, on some set $\mathcal{Z}^{\prime}$, such that
$\mathbb{E}_{z^{\prime}\sim Z^{\prime}}h(\mathbb{P}\left(Y|Z^{\prime}=z^{\prime}\right))\leq\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\text{ and }\left\|\mu^{\prime}_{0}-\mu^{\prime}_{1}\right\|_{TV}=\left\|\mu_{0}-\mu_{1}\right\|_{TV}.$ (7)
Here, similarly to $\mu_{a}$, $\mu^{\prime}_{a}$ are the distributions induced by $T^{\prime}$ on $\mathcal{Z}^{\prime}$. In words, for every representation, we can find an invertible representation of the same data which satisfies at least as good a fairness constraint, and has at least as good performance as the original. In particular, this implies that when one searches for optimal performance representations, it suffices to only search among the invertible ones.
The proof proceeds by observing that if an atom $z\in\mathcal{Z}$ has more than one parent on the same side (i.e. $S_{0}$ or $S_{1}$), then one can split this atom into two, with each having less parents. However, the details of this construction are somewhat intricate. The full argument can be found in Section C.
@@ 5 The Model Independent Optimization Problem
This Section describes the specific numerical problem, MIFPO, that is used to compute the Fairness-Performance Pareto Front. Section 5.1 discusses the factorisation of representations, in Section 5.2 the MIFPO problem is defined, and in Section 5.3 we describe how MIFPO is instantiated from data.
To motivate MIFPO, let us write again some of the terms involved in the expressions of the performance cost and the fairness constraints $E$ and $F$, as defined in Section 3. To simplify the exposition, we use the discrete summation notation, but similar expressions can also be written using integration and densities. We have
$\mathbb{P}\left(z\right)=\sum_{x,a}\mathbb{P}\left(z|x,a\right)\mathbb{P}\left(x,a\right)\text{ and }\mathbb{P}\left(z|a\right)=\sum_{x}\mathbb{P}\left(z|a,x\right)\mathbb{P}\left(x,a\right)/\mathbb{P}\left(a\right).$ (8)
Moreover,
$\mathbb{P}\left(Y|z\right)=\sum_{x,a}\mathbb{P}\left(Y|x,a,z\right)\mathbb{P}\left(a,x|z\right)=\sum_{x,a}\mathbb{P}\left(Y|x,a\right)\mathbb{P}\left(z|a,x\right)\mathbb{P}\left(a,x\right)/\mathbb{P}\left(z\right),$ (9)
where we have used $\mathbb{P}\left(Y|x,a,z\right)=\mathbb{P}\left(Y|x,a\right)$, which holds due to the independence condition (1) of the representation. We observe that one can write the performance cost and the fairness constraint, $E$ and $F$, in terms of $\mathbb{P}\left(a\right),\mathbb{P}\left(x,a\right),\mathbb{P}\left(Y|x,a\right)$, and of the representation $\mathbb{P}\left(z|x,a\right)$. We refer to these quantities as the basic probabilities. It follows that any two problems having the same basic probabilities, will have the same $E$ and $F$.
One way to view the construction of the MIPFO problem is simply as a discretization, where we model the above basic probabilities in an abstract way, on a finite set, without the need to consider the details of how the representation $\mathbb{P}\left(z|x,a\right)$ is implemented. However, discretization of the features $x\in\mathbb{R}^{d}$ directly would likely always be too crude. In the following Section we show that in fact one only needs to discretize a much smaller space, namely $\Delta_{\mathcal{Y}}$ the set of distributions over $Y$ values, which is typically easy to achieve. In Section 5.3 we show how factorisation is used in practice.
Figure 2: (a) The MIFPO Setting (b) Factorisation Through $\Delta_{\mathcal{Y}}$
5.1 Factorisation
The general idea behind factorisation is the observation that in a given representation, individual source points $x,x^{\prime}$ may be merged if they have the same $\rho_{x}=\mathbb{P}\left(Y|x\right)$ distribution, without affecting the measurements $E$ and $F$. We now carry out this argument formally, with Figure 2(b) illustrating the construction. For clarity, we use the notation of Section 4 for finite sets. The extension to the continuous case is discussed in the end of the Section.
Suppose that we have a representation on some, possibly very large source space $S=S_{0}\cup S_{1}$, with representation space $\mathcal{Z}$. Introduce a map $\Psi:S_{0}\rightarrow\Delta_{\mathcal{Y}}$, given by $\Psi(v)=\rho_{v}$. Then we can define a distribution $\bar{\beta}_{0}$ on $\Delta_{\mathcal{Y}}$, and the transition kernel $\bar{T}_{0}:\mathcal{Z}\times\Delta_{\mathcal{Y}}\rightarrow[0,1]$ by
$\bar{\beta}_{0}(\rho)=\sum_{v\in\Psi^{-1}(\left\{\rho\right\})}\beta_{0}(v)\hskip 5.69054pt\text{ and }\bar{T}_{0}(z,\rho)=\sum_{v\in\Psi^{-1}(\left\{\rho\right\})}\frac{\beta_{0}(v)}{\bar{\beta}_{0}(\rho)}T_{0}(z,v),\text{ for all $\rho\in\Delta_{\mathcal{Y}}$}.$ (10)
Similar construction can be made for $a=1$, defining $\bar{\beta}_{1}$, $\bar{T}_{1}$. Observe that together, $\bar{\beta}_{a}$ and $\bar{T}_{a}$ define a representation on $\Delta_{\mathcal{Y}}$ with values in $\mathcal{Z}$ (technically, the source space is $\Delta_{\mathcal{Y}}^{0}\cup\Delta_{\mathcal{Y}}^{1}$, where $\Delta_{\mathcal{Y}}^{a}$ are two disjoint copies of $\Delta_{\mathcal{Y}}$). Note that it is understood here that the distribution of $Y$ at point $\rho\in\Delta_{\mathcal{Y}}$ is taken to be $\rho$ itself. The crucial property of this new representation is that the induced distributions on $\mathcal{Z}$ coincide with those of the old one. That is,
$\mu_{a}(z)=\bar{\mu_{a}}(z)\text{ and }\mathbb{P}_{T}\left(Y|Z=z\right)=\mathbb{P}_{\bar{T}}\left(Y|Z=z\right)$ (11)
for all $z\in\mathcal{Z}$. This property follows simply by substituting the definitions, although the derivation is slightly notationally heavy. The full details may be found in Supplementary Section F.2.
The meaning of (11) is that from $\mathcal{Z}$ one can not distinguish between the representations $T$ and $\bar{T}$, and in particular they have the same $E$ and $F$. We say that $T$ is factorised through $\Delta_{\mathcal{Y}}$ since it can be represented as a composition, of first applying $\Psi$ and then applying $\bar{T}$.
When the feature space $S$ is continuous, e.x. $\mathbb{R}^{d}$, the construction of $\bar{T}$ is similar. Indeed, $\Psi$ can be defined in the same way, $\bar{\beta}_{a}$ are defined as push-forwards $\bar{\beta}_{a}=\Psi\beta_{a}$, and $\bar{T}_{a}(z,\rho)$ is defined as conditional expectation, $\bar{T}_{a}(z,\rho)=\mathbb{E}_{v\sim\beta_{0}}T_{a}(z,v)|\left(\Psi(v)=\rho\right)$.
@@ 5.2 MIFPO Definition
Fix an integer $k\geq 2$. The representation space $\mathcal{Z}$ will be a finite set which can be written as $\mathcal{Z}=S_{0}\times S_{1}\times[k]$, where $[k]:=\left\{1,2,\ldots,k\right\}$. That is, every point $z\in\mathcal{Z}$ corresponds to some triplet $(u,v,j)$, with $u\in S_{0},v\in S_{1},j\in[k]$. The motivation for this choice of $\mathcal{Z}$ is as follows: First, note the by the invertibility Theorem 4.1, in an optimal representation, every $z\in\mathcal{Z}$ is associated with a pair of parents, $(u,v)$. Moreover, it can be shown that for every pair $(u,v)$, it is sufficient to have at most $k$ points $z$ which have $(u,v)$ as parents, to approximate all representations on all possible spaces $\mathcal{Z}^{\prime}$, in terms of their $E$ and $F$ quantities, to a given degree $\varepsilon$. Here, $k$ would depend on $\varepsilon$ and on the function $h$, but not on the particular representations, or even the sizes $\left|S_{0}\right|,\left|S_{1}\right|$, etc. See Section D in Supplementary Material, for additional details. As a result, choosing $\mathcal{Z}=S_{0}\times S_{1}\times[k]$ should be sufficient to model all representations well. We have used $k=10$ in all experiments.
Next, the variables of the problem, modeling the representation itself, will be denoted by $r_{u,v,j}$ and $r_{v,u,j}$, for $(u,v,j)\in\mathcal{Z}$. These values correspond to the probabilities $r_{u,v,j}=\mathbb{P}\left(Z=(u,v,j)|(X,A)=(u,0)\right)$ and $r_{v,u,j}=\mathbb{P}\left(Z=(u,v,j)|(X,A)=(v,1)\right)$. Note that the order of $u,v$ in $r_{u,v,j}$ and $r_{v,u,j}$ determines whether the weight is transferred from $S_{0}$ or $S_{1}$. The values $r$ correspond to the values of $T_{0}(z,u),T_{1}(z,v)$ in Section 4. However, note that due to the special structure of the representation here, many values will be 00, such as for instance $T_{0}((u,v,j),u^{\prime})$, whenever $u\neq u^{\prime}$. The $r$ notation accounts for this, and is thus clearer. The whole setup is illustrated in Figure 2(a).
@@
$r_{u,v,j}\geq 0,\hskip 5.69054ptr_{v,u,j}\geq 0\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z}$ (12)
$\sum_{v\in S_{1},j\in[k]}r_{u,v,j}=1\hskip 5.69054pt\forall u\in S_{0}\text{ and }\sum_{u\in S_{0},j\in[k]}r_{v,u,j}=1\hskip 5.69054pt\forall v\in S_{1}.$ (13)
Recall from Section 4 that the additional parameters of the problem are the conditional distributions $\beta_{0}(u),\beta_{1}(v)$, on $S_{0}$ and $S_{1}$, modeling $\mathbb{P}\left((X,A)|A=0\right)$ and $\mathbb{P}\left((X,A)|A=1\right)$ respectively. The quantities $\alpha_{a}=\mathbb{P}\left(A=a\right)$, and the conditional $Y$ distributions, $\rho_{u},\rho_{v}\in\Delta_{\mathcal{Y}}$, modelling $\rho_{u}=\mathbb{P}\left(Y=\cdot|(X,A)=(u,0)\right)$, $\rho_{v}=\mathbb{P}\left(Y=\cdot|(X,A)=(v,1)\right)$. Here we assume that these quantities are given, and in Section 5.3 below we describe how one may estimate them from the data.
With these preparations, we are ready to write the performance cost (5) in the new notation:
$E_{r}=\sum_{z=(u,v,j)}\left(\alpha_{0}\beta_{0}(u)r_{u,v,j}+\alpha_{1}\beta_{1}(v)r_{v,u,j}\right)h\left(\frac{\rho_{u}\alpha_{0}\beta_{0}(u)r_{u,v,j}+\rho_{v}\alpha_{1}\beta_{1}(v)r_{v,u,j}}{\alpha_{0}\beta_{0}(u)r_{u,v,j}+\alpha_{1}\beta_{1}(v)r_{v,u,j}}\right).$ (14)
Indeed, observe that due to the structure of our representations, every $z$ has two parents, and we have $\mathbb{P}\left(Z=z\right)=\left(\alpha_{0}\beta_{0}(u)r_{u,v,j}+\alpha_{1}\beta_{1}(v)r_{v,u,j}\right)$. Similarly, $\mathbb{P}\left(Y|Z=z\right)$ is computed via (9) and substituted inside $h$ to obtain (14). As we show in Supplementary E, the cost (14) is a concave function of the parameters $r$.
We now proceed to discuss the fairness constraint $F$. Recall that we define $\mu_{a}(z)=\mathbb{P}\left(Z=z|A=a\right)$, for $a\in\left\{0,1\right\}$. For $z=(u,v,j)$ we have then $\mu_{0}((u,v,j))=\beta_{0}(u)r_{u,v,j}$ and $\mu_{1}((u,v,j))=\beta_{1}(v)r_{v,u,j}$. We thus can write
$F_{r}=\left\|\mu_{0}-\mu_{1}\right\|_{TV}=\frac{1}{2}\sum_{z}\left|\mu_{0}(z)-\mu_{1}(z)\right|=\frac{1}{2}\sum_{(u,v,j)}\left|\beta_{0}(u)r_{u,v,j}-\beta_{1}(v)r_{v,u,j}\right|$ (15)
and the Fairness constraint, for a given $\gamma\in[0,1]$ is simply
$F_{r}\leq\gamma.$ (16)
Although the constraint (16) is convex in the variables $r$, and can be incorporated directly into most optimisation frameworks, it may be significantly more convenient to work with equality constraints. Using the following Lemma, we can find equivalent equality constraints in a particularly simple form.
Lemma 5.1.
As consequence, if we find distributions $\phi_{0},\phi_{1}\in\Delta_{\mathcal{Z}}$ such that $\mu_{0}+\gamma\phi_{0}=\mu_{1}+\gamma\phi_{1}$ holds, then we know that (16) also holds, and conversely, if (16) holds, then distributions as above exist.
Using this observation, we introduce new variables, $\phi_{u,v,j}$ and $\phi_{v,u,z}$, for every $(u,v,j)\in\mathcal{Z}$, which correspond to $\phi_{0}((u,v,j))$ and $\phi_{1}((u,v,j))$ respectively. These variables will be required to satisfy the following constraints:
$r_{u,v,j}\geq 0,\hskip 5.69054ptr_{v,u,j}\geq 0\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z}$ (17)
$\sum_{u,v,j}r_{u,v,j}=1\text{ and }\sum_{u,v,j}r_{v,u,j}=1$ (18)
$\beta_{0}(u)r_{u,v,j}+\gamma\phi_{u,v,j}=\beta_{1}(v)r_{v,u,j}+\gamma\phi_{v,u,j}\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z}.$ (19)
@@
For a fixed finite ground set $S=S_{0}\cup S_{1}$, the problem parameters are the weight $\alpha_{0}$, the distributions $\beta_{0},\beta_{1}$, on $S_{0}$ and $S_{1}$ respectively, and the distributions $\rho_{x}\in\Delta_{\mathcal{Y}}$ for every $x\in S$. The problem variables are $\left\{r_{u,v,j},r_{v,u,j},\phi_{u,v,j},\phi_{v,u,j}\right\}_{(u,v,j)\in\mathcal{Z}}$ as defined above. We are interested in minimising the concave function (14), subject to the constraints (12)-(13) and (17)-(19).
@@
The relationship between MIFPO and the Optimal Transport problem is detailed in Section B.
5.3 Constructing a MIFPO Instance
In this Section we discuss the estimation of the quantities $\alpha$,$\beta$ and $\rho$ from the data, and specify how the data is used to construct the MIFPO problem instance.
For clarity we treat the case of binary $Y$, which is also the case for all datasets in Balunović et al., 2022a , on which we base our experiments. Note that for binary $Y$, $\Delta_{\mathcal{Y}}$ is simply the interval $[0,1]$, with $p\in[0,1]$ corresponding to the probability $p=\mathbb{P}\left(Y=1|x,a\right)$.
The computation proceeds in 3 steps, schematically shown as Algorithm 1. The input is a labeled dataset $\left\{(x_{i},a_{i},y_{i})\right\}_{i\leq N}$, with $x_{i}\in\mathbb{R}^{d}$, $a_{i},y_{i}\in\left\{0,1\right\}$, and two integers, $L,k$, which will specify the approximation degrees for the discretisation and for the approximation space, respectively. In Step 1., separately for the parts of the data corresponding to $A=0$ and $A=1$, we learn calibrated classifiers $c_{0},c_{1}:\mathbb{R}^{d}\rightarrow[0,1]$, which represent the probabilities $c_{a}(x)=\mathbb{P}\left(Y=1|X=x,A=a\right)=\rho_{a}(x)$. Note the following points: (a) In most applications, the true values of $\mathbb{P}\left(Y=1|X=x,A=a\right)$ are non-binary (i.e. $\neq 0,1$), and thus the $\left\{0,1\right\}$-valued labels are only noisy observations of the actual probabilities. Thus, some sort of estimator must be used to learn the probabilities. (b) Typical classification objectives tend to yield in an imprecise estimation of $\mathbb{P}\left(Y=1|X=x,A=a\right)$. However, this can usually be fixed by the process of calibration in post-processing, see Niculescu-Mizil and Caruana, (2005), Kumar et al., (2019), (Berta et al.,, 2024). In Step 2. we define the sets $H_{a}=\left\{c_{a}(x)\hskip 5.69054pt|\hskip 5.69054ptx\in X_{a}\right\}\subset[0,1]$, and approximate them by histograms with $L$ bins. That is, for each $a$, we find $L$ points $\rho_{a}^{1},\ldots,\rho_{a}^{L}\in[0,1]$, and weights $\beta_{a}^{1},\ldots,\beta_{a}^{L}$, such that the distribution with centers $\rho_{a}^{i}$ and weights $\beta_{a}^{i}$ approximates $H_{a}$ well. We typically take uniform bins, i.e. simply $\rho_{a}^{l}=\frac{l-1}{L-1}$, for $l=1,\ldots,L$. See Figure 3(b) for an example of such histograms on real data. Finally, in Step 3. we solve the MIFPO problem, as given in Definition 5.1. The ground set here is $S=S_{0}\cup S_{1}$, with $S_{0}=\left\{1,\ldots,L\right\}$ and $S_{1}=\left\{1^{\prime},\ldots,L^{\prime}\right\}$, where we have added in $S_{1}$ to signify that the sets are disjoint copies in $S$, and $\left|S\right|=2L$. The problem parameters are $\alpha_{a}^{l},\beta_{a}^{l},\rho_{a}^{l}$, as computed in the previous step, and the parameter $k$ determining the size of $\mathcal{Z}$.

### v2

@@ 3.3 The Invertibility Theorem
In this section we define the notion of invertibility for representations, and show that considering invertible representations is sufficient for computing the Pareto front.
@@
A representation $Z$ on a set $\mathcal{Z}$ is invertible if for every $z\in\mathcal{Z}$ and every $a\in\left\{0,1\right\}$, there is at most one $x\in\mathcal{X}$ such that $\mathbb{P}\left(Z=z|X=x,A=a\right)>0$. In words, a representation is invertible, if any given $z$ can be produced by at most two original features $(x,a)$, and at most one for each value $a$. For $z\in\mathcal{Z}$, we say that an $(x,a)$ is a parent of $z$ if $\mathbb{P}\left(Z=z|X=x,A=a\right)>0$.
@@
Let $Z$ be any representation of $(X,A)$ on a set $\mathcal{Z}$. Then there exists an invertible representation $Z^{\prime}$ of $(X,A)$, on some set $\mathcal{Z}^{\prime}$, such that
$\mathbb{E}_{z^{\prime}\sim Z^{\prime}}h(\mathbb{P}\left(Y|Z^{\prime}=z^{\prime}\right))\leq\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\text{ and }D_{TV}(Z^{\prime})=D_{TV}(Z).$ (10)
In words, for every representation, we can find an invertible representation of the same data which satisfies at least as good a fairness constraint, and has at least as good performance as the original. In particular, this implies that when one searches for optimal performance representations, it suffices to only search among the invertible ones.
The proof proceeds by observing that if an atom $z\in\mathcal{Z}$ has more than one parent for a fixed $a$, then one can split this atom into two, with each having less parents. However, the details of this construction are somewhat intricate and the full argument can be found in Section C.
Although in this paper we concentrate on the case of binary sensitive variable, we note that Theorem 3.1 may be extended to multi valued attributes, with a similar argument. In that case, invertibility would mean that every $z\in\mathcal{Z}$ would still have at most two parents, $u,v$, corresponding to different values $a,a^{\prime}$ of $A$.
@@ 4 The Model Independent Optimization Problem
In this Section we motivate and introduce the MIFPO optimisation problem, and then discuss the full Pareto front computation procedure starting from the raw data.
Figure 2: (a) The MIFPO Setting (b) Distribution of $P(Y=1|X,A)$ for each group across datasets.
@@ 4.1 MIFPO Definition
For the purposes of this Section, we assume that the $x$ feature space $\mathcal{X}$ is finite. In the next section, Section 4.2, we describe how we obtain such finite spaces by using the factorization result and discretizing $\Delta_{\mathcal{Y}}$. Note, however, that the full original, possibly high dimensional feature space $\mathcal{X}$, is never discretized.
Write $S_{0}=\left\{(x,0)\hskip 5.69054pt|\hskip 5.69054ptx\in\mathcal{X}\right\}=\mathcal{X}\times\left\{0\right\}$, and similarly $S_{1}=\mathcal{X}\times\left\{1\right\}$, for the two halves of the full feature space, $\mathcal{X}\times\mathcal{A}=S_{0}\cup S_{1}$.
Parameters: The MIFPO parameters model the data distribution and are as follows: (a) the probability distributions $\beta_{0}\in\Delta_{S_{0}}$ and $\beta_{1}\in\Delta_{S_{1}}$, on $S_{0}$ and $S_{1}$ respectively, modeling $\mathbb{P}\left((X,A)|A=0\right)$ and $\mathbb{P}\left((X,A)|A=1\right)$ respectively, i.e. the distribution of the data features on each sensitive subgroup. (b) The subgroup proportions $\alpha_{a}=\mathbb{P}\left(A=a\right)$, and (c) the conditional $Y$ distributions, $\rho_{u},\rho_{v}\in\Delta_{\mathcal{Y}}$, modeling $\rho_{u}=\mathbb{P}\left(Y=\cdot|(X,A)=u\right)$ when $a=0$ or $\rho_{v}=\mathbb{P}\left(Y=\cdot|(X,A)=v\right)$ when $a=1$.
Representation Space: Perhaps the first question one can ask when constructing a representation of the data as above is: How large the representation space should be? We now answer this question using the theory of Section 3.
Fix an integer $k\geq 2$. The representation space $\mathcal{Z}$ will be a finite set which can be written as
$\mathcal{Z}=S_{0}\times S_{1}\times[k],$ (11)
where $k]:=\left\{1,2,\ldots,k\right\}$. That is, every point $z\in\mathcal{Z}$ corresponds to some triplet $(u,v,j)$, with $u\in S_{0},v\in S_{1},j\in[k]$. To explain this choice, recall that by the Invertibility result, we know that we may consider only invertible representations. In such representations, every point $z\in\mathcal{Z}$ is indexed by a pair of parents $(u,v)\in S_{0}\times S_{1}$, suggesting that we may index the points by $S_{0}\times S_{1}$ to begin with. Next, for a given such pair $(u,v)$, we may ask how many points $z$ should have the same pair $(u,v)$ as their parents? In Supplementary Section D, we show that using $k$ points for every pair, we can obtain uniform approximation over all representations. That is, given a degree of approximation $\varepsilon$, Lemma D.1 provides a bound on $k$ which is sufficient to obtain such approximation. While such a bound would clearly depend on $\varepsilon$, we not that it does note depend on the sizes $\left|S_{0}\right|,\left|S_{1}\right|$. These considerations explain the choice of (11) as the representation space. We have used $k=5$ in all experiments.
Variables: The variables of the problem model the representation itself. They will be denoted by $r_{u,v,j}^{a}$ for $(u,v,j)\in\mathcal{Z}$ and $a\in\mathcal{A}$, and model the probabilities $r_{u,v,j}^{a}=\mathbb{P}\left(Z=(u,v,j)|(X,A)=s\right)$, where either $a=0$ and $s=u\in S_{0}$, or $a=1$ and $s=v\in S_{1}$ for some $v\in\mathcal{X}$. That is, for $a=0$, points $u$ transition to $(u,v,j)$ for some $v\in S_{1},j\in[k]$, and similarly for $a=1$, points $v$ transition to $(u,v,j)$ for some $u\in\mathcal{X},j\in[k]$. This notation preserves our convention that $(u,v,j)\in Z$ has $u$ and $v$ as its only parents. The situation is illustrated in Figure 2(a).
@@
$r_{u,v,j}^{a}\geq 0,\hskip 5.69054pt\forall(u,v,j)\in\mathcal{Z},\forall a\in\mathcal{A}$ (12)
$\sum_{v\in S_{1},j\in[k]}r_{u,v,j}^{0}=1\hskip 5.69054pt\forall u\in S_{0}\text{ and }\sum_{u\in S_{0},j\in[k]}r_{u,v,j}^{1}=1\hskip 5.69054pt\forall v\in S_{1}$ (13)
Performance Objective and Fairness Constraints: With these preparations, we are ready to write the performance cost (5) in the new notation:
$\begin{split}E_{r}=\sum_{z=(u,v,j)}\left[\alpha_{0}\beta_{0}(u)r_{u,v,j}^{0}+\alpha_{1}\beta_{1}(v)r_{u,v,j}^{1}\right]\cdot h\left(\frac{\rho_{u}\alpha_{0}\beta_{0}(u)r_{u,v,j}^{0}+\rho_{v}\alpha_{1}\beta_{1}(v)r_{u,v,j}^{1}}{\alpha_{0}\beta_{0}(u)r_{u,v,j}^{0}+\alpha_{1}\beta_{1}(v)r_{u,v,j}^{1}}\right).\end{split}$ (14)
Indeed, observe that due to the structure of our representations, every $z$ has two parents, and we have $\mathbb{P}\left(Z=z\right)=\left(\alpha_{0}\beta_{0}(u)r_{u,v,j}^{0}+\alpha_{1}\beta_{1}(v)r_{u,v,j}^{1}\right)$. Similarly, $\mathbb{P}\left(Y|Z=z\right)$ is computed via
$\mathbb{P}\left(Y|z\right)=\sum_{x,a}\mathbb{P}\left(Y|x,a,z\right)\mathbb{P}\left(a,x|z\right)=\sum_{x,a}\mathbb{P}\left(Y|x,a\right)\mathbb{P}\left(z|a,x\right)\mathbb{P}\left(a,x\right)/\mathbb{P}\left(z\right),$
and substituted inside $h$ to obtain (14). As we show in Supplementary E, the cost (14) is a concave function of the variables $r$.
We now proceed to discuss the fairness constraint. Recall that we define $\mu_{a}(z)=\mathbb{P}\left(Z=z|A=a\right)$, for $a\in\left\{0,1\right\}$. For $z=(u,v,j)$ we have then $\mu_{a}((u,v,j))=\beta_{a}(u)r_{u,v,j}^{a}$, for $a\in\left\{0,1\right\}$, and we can write
$\begin{split}D_{TV}(Z)=\left\|\mu_{0}-\mu_{1}\right\|_{TV}=\frac{1}{2}\sum_{z}\left|\mu_{0}(z)-\mu_{1}(z)\right|=\frac{1}{2}\sum_{(u,v,j)}\left|\beta_{0}(u)r_{u,v,j}^{0}-\beta_{1}(v)r_{u,v,j}^{1}\right|\end{split}$ (15)
and the Fairness constraint, for a given $\gamma\in[0,1]$, is thus simply
$\frac{1}{2}\sum_{(u,v,j)}\left|\beta_{0}(u)r_{u,v,j}^{0}-\beta_{1}(v)r_{u,v,j}^{1}\right|\leq\gamma.$ (16)
@@
For a fixed finite ground set $\mathcal{X}\times\mathcal{A}=S_{0}\cup S_{1}$, the problem parameters are the weight $\alpha_{0}$, the distributions $\beta_{0},\beta_{1}$, on $S_{0}$ and $S_{1}$ respectively, and the distributions $\rho_{x}\in\Delta_{\mathcal{Y}}$ for every $x\in S_{0}\cup S_{1}$. The problem variables are $\left\{r_{u,v,j}^{0},r_{u,v,j}^{1}\right\}_{(u,v,j)\in\mathcal{Z}}$ as defined above. We are interested in minimizing the concave function (14), subject to the constraints (12), (13), and (16).
@@
The relationship between MIFPO and the Optimal Transport problem is detailed in Supplementary B.
Finally, observe that the MIFPO constraints above linear, with (16) being equivalent to two linear inequality constraints. We note that these constraints may be replace by equivalent linear equality constraints, via appropriate slack variables, which is more convenient in practice. See Supplementary F for details.
4.2 The Full Algorithm
In this Section we summarize the full Pareto front computation algorithm, including the estimation of the MIFPO parameters $\alpha$$\beta$ and $\rho$ as discussed above.
Let $D=\left\{((x_{i},a_{i}),y_{i})\right\}_{i\leq N}$ be the dataset, and write $D_{a}=\left\{((x_{i},a_{i}),y_{i})\in D\hskip 5.69054pt|\hskip 5.69054pta_{i}=a\right\}$, so that $D=D_{0}\cup D_{1}$.
The algorithm proceeds in the following steps: Step 1: we learn the probability estimators $c_{0},c_{1}:\mathbb{R}^{d}\rightarrow\Delta_{\mathcal{Y}}$, separately on $D_{0}$ and $D_{1}$. These estimators should approximate the optimal Bayes classifier (Section 3.2). Note that such estimation of probabilities is well studied, and is known as calibration, see Niculescu-Mizil and Caruana, (2005), Kumar et al., (2019), (Berta et al.,, 2024).
Step 2: (Discretization and Parameter estimation) For a given integer $L>0$, the space $\Delta_{\mathcal{Y}}$ is discretized into $L$ bins. This corresponds to taking the ground sets $S_{0},S_{1}$ in MIFPO to be of size $L$. The data $D_{a}$ is then mapped into the $L$ bins using $c_{a}$. The distribution $\beta_{a}(w)$ is then simply measures the proportion of points $\left\{c_{a}(x)\right\}_{(x,a)\in D_{a}}$ that fall into bin $w\leq L$. Finally, for every bin $w$, we choose an arbitrary point inside that bin as the representative distribution, $\rho_{w}$. The parameters $\alpha_{a}$ are estimated simply by $\alpha_{a}=\left|D_{a}\right|/\left|D\right|$. Note that for binary $Y$, $\Delta_{\mathcal{Y}}$ is simply the interval $[0,1]$, which is trivial to discretize. See Figure 2(b) for an example of such histograms on real data. We note that one could easily consider more complex discretization schemes, such as clustering, which could be applied efficiently to multi label problems. See Supplementary J for a discussion.
Step 3: For a given a fairness threshold $\gamma>0$, we can now construct the MIFPO instance, Definition 4.1, with $\left|S_{0}\right|=\left|S_{1}\right|=L$, the additional approximation parameter $k$, and $\alpha,\beta,\rho$ as discussed above. As discussed in Section 4.1, we found it sufficient to use $k=5$ throughout the paper. The MIFPO is then solved using the existing methods, as detailed in Section 5. The full algorithm is schematically show as Algorithm 1, Supplementary H.1.